OpenStudy (anonymous):

Please help, will fan and medal! The length and width of a rectangle are consecutive even integers. The area of the rectangle is 48 square units. What are the length and width of the rectangle. Separate the answers with a comma

OpenStudy (anonymous):

width * length = 48. Looking for the factors of 48 you get: 2, 2, 2, 2, 3 (2*2*2*2*3=48). Now, you just need to figure out how to multiply up these numbers to get two consecutive even integers. The answer would be 2*2*2=8 and 2*3=6. 6, 8

OpenStudy (anonymous):

Thank you :)